OpenStudy (anonymous):

Are 2 times x plus 4 all over 3 times x plus 6 and 2 over 3 equivalent expressions? Using complete sentences explain why or why not.

OpenStudy (anonymous):

Yes. The left side simplifies to equal the right side no matter what.

OpenStudy (mikezack123):

omg that answer is so wrong @xchessmachinex

OpenStudy (aravindg):

I don't think so? It is right.

OpenStudy (mathmale):

Would one of you please write out this statement mathematically? It's hard to decipher the problem when expressed only in words.

OpenStudy (mathmale):

@MikeZack123 : I invite you to do that.

OpenStudy (mikezack123):

@mathmale I decline your invitation.

OpenStudy (mathmale):

that's ok, but you did seem to be absolutely convincned that the other person was wrong.

OpenStudy (mikezack123):

dont h8 plz

OpenStudy (aravindg):

Next time think twice before you make a report like that.

OpenStudy (mikezack123):

@AravindG omg i said don't h8 plz omg ur h8ing on me omg

OpenStudy (aravindg):

Nope just giving you some advice :) Do you realize why his answer is right?

OpenStudy (mikezack123):

its not srry not srry

OpenStudy (aravindg):

What do you mean?

OpenStudy (mikezack123):

its rong okai? don h8 plz

OpenStudy (aravindg):

You mean you realize you are wrong?

OpenStudy (mikezack123):

no srry u are wrong, okai? plz don h8. #stoph8ing2014

OpenStudy (aravindg):

\[\dfrac{2x+4}{3x+6}=\dfrac{2}{3}\dfrac{(x+2)}{(x+2)}=\dfrac{2}{3}\]

OpenStudy (mikezack123):

wat are that?

OpenStudy (aravindg):

I showed you how we get 2/3.

OpenStudy (mikezack123):

i are wrong srry i am dumb :( thnx fer ur help :)

OpenStudy (aravindg):

yw :) (no h8 at all!)

OpenStudy (mikezack123):

yes no h8 :) thnx again